# Towards Reliable Evaluation of Adversarial Robustness for Spiking Neural Networks

### v1

@@ 1 Introduction
Spiking Neural Networks (SNNs) have emerged as brain-inspired models that emulate the sparse, event-driven processing of biological neural systems [yin2021accurate, wu2018spatio, cai2025robust, yan2025efficient, fang2023parallel]. By transmitting information through discrete spikes instead of continuous activations, SNNs achieve high energy efficiency and rich temporal dynamics, making them attractive for neuromorphic and low-power applications [pei2019towards, roy2019towards, davies2018loihi, orchard2021efficient, ma2024darwin3]. However, the discontinuous and binary nature of spike activations hinders the application of gradient-based adversarial attacks, making reliable robustness evaluation difficult [athalye2018obfuscated].
To overcome the non-differentiability of spike activations, surrogate gradient methods approximate the spike derivative with a smooth function [wu2018spatio, mukhoty2023direct]. While this enables backpropagation, it introduces approximation errors that distort gradients, especially under adversarial perturbations where accurate input gradients are essential [athalye2018obfuscated, athalye2018synthesizing]. As a result, gradient-based robustness evaluation in SNNs remains unreliable.
Several studies have refined surrogate gradients [bu2023rate, hao2023threaten, lun2025towards, xu2025attacking], yet existing methods lack input-dependent designs and theoretical analysis of gradient vanishing. Moreover, most works overlook the development of constrained optimization algorithms for adversarial attacks, which are equally crucial for reliable robustness evaluation.
@@
  • We design a new adversarial optimization algorithm with adaptive step-size control, which achieves more stable and efficient convergence even when gradients are imprecise.
@@
  • Our findings reveal that the adversarial robustness of current SNNs has been significantly overestimated, highlighting the urgent need to rethink how to efficiently generate reliable adversarial samples for training more genuinely robust SNN models.
@@ 2.1 Gradient Approximation Methods in SNNs
Recent studies have shown that SNNs are vulnerable to adversarial perturbations [ding2022snn, liang2021exploring]. Surrogate gradient methods approximate the spike activation with a continuous function to enable Spatial-Temporal Backpropagation (STBP) [wu2018spatio]. The Potential-Dependent Surrogate Gradient (PDSG) method further incorporates the variance of membrane potential distributions into the surrogate computation [lun2025towards]. However, PDSG derives the surrogate function from batch-level statistics rather than individual samples, preventing adaptive modeling of sample-specific dynamics and lacking theoretical justification for its formulation.
Another line of research estimates input gradients by differentiating the SNN loss with respect to neurons’ average firing rates. Representative methods include Backward Pass Through Rate (BPTR) [ding2022snn] and Rate Gradient Approximation (RGA) [bu2023rate]. The Hybrid Adversarial attack based on Rate and Temporal information (HART) [hao2023threaten] jointly considers rate- and time-based dependencies, but its derivation is limited to a specific neuronal model and lacks theoretical validation for generalization to broader neuron types.
@@ 2.2 Adversarial Defense Methods for SNNs
Early studies adapted adversarial training for SNNs, such as adversarial input augmentation and min–max optimization [madry2017towards]. Later approaches, including Regularized Adversarial Training (RAT), constrain the Lipschitz constant of each layer via weight regularization to stabilize learning [ding2022snn], while Sparsity Regularization (SR) introduces a sparse gradient penalty into the loss to enhance robustness when combined with adversarial training [liu2024enhancing].
More recent works have strengthened SNN adversarial defense from an architectural perspective [ding2024enhancing, xu2024feel]. For instance, synaptic gating mechanisms dynamically regulate spike transmission to suppress adversarial noise propagation [ding2024enhancing], and frequency-domain encoding schemes improve robustness by representing spiking information in more stable spectral domains [xu2024feel].
Figure 1: (a). Curves of the surrogate gradient function $g(x)$ and its integral under different sharpness parameters $\alpha$. (b). Illustration of the adaptive mechanism of the surrogate gradient sharpness parameter in ASSG during adversarial attacks.
@@ 3.1 Adversarial Attack and Training
Adversarial attacks can be viewed as a constrained optimization problem in the input space of a neural network [moosavi2016deepfool, carlini2017towards, madry2017towards, croce2020reliable, dong2018boosting]. For a classification model $\boldsymbol{f}_{\boldsymbol{\theta}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$, let $\boldsymbol{x}\in\mathbb{R}^{d}$ denote a benign input sample with the ground-truth label $y$. The objective of an untargeted adversarial attack is to find a perturbed input $\boldsymbol{x^{\prime}}$ that maximizes the classification loss $\mathcal{L}$ while remaining within a small $L_{p}$-norm ball around the original input:
@@
here, $\varepsilon$ defines the perturbation bound that controls the perceptual similarity between $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$. In practice, gradient-based optimization methods such as Projected Gradient Descent (PGD) and Auto-PGD (APGD) are commonly used to approximate the solution of this problem under various norm constraints (e.g. $L_{\infty}$, $L_{2}$[madry2017towards, croce2020reliable].
Adversarial Training (AT) is one of the most fundamental methods for improving a network’s adversarial robustness, and it has also been widely adopted in SNN-related studies [ding2022snn, ding2024robust, ding2024enhancing, xu2024feel, liu2024enhancing, mukhoty2025improving, kundu2021hire]. The objective of AT can be formulated as a min–max optimization problem:
@@
where $\mathcal{D}$ denotes the data distribution. TRadeoff-inspired Adversarial DEfense via Surrogate-loss minimization (TRADES) is one of the most influential variants of AT [zhang2019theoretically, pang2022robustness], which is also employed as a training framework for developing robust SNNs in this work.
@@ 3.2 Spiking Neurons and Surrogate Gradients
To evaluate the generality of the proposed attack, we employ several neuron models in addition to the standard LIF neuron, including its variant LIF-2, the Integrate-and-Fire (IF) neuron, and the Parallel Spiking Neuron (PSN) [fang2023parallel]. Detailed configurations of these models are provided in the Supplementary Materials.
@@
where $\alpha$ controls the sharpness of the surrogate gradient.
@@ 4.1 Theoretical Analysis on Gradient Vanishing
$H(x)$ is discontinuous and has zero derivatives everywhere in its continuous domain. Therefore, surrogate gradient is employed to approximate a non-zero and smooth derivative for training or adversarial attack. We define the surrogate gradient as the following general class of functions:
@@
Unlike the definition given by Mukhoty et al. [mukhoty2023direct], we impose the normalization constraint $\int_{-\infty}^{\infty}g(x)\,dx=1$. This property guarantees that the integral of $g(x)$ behaves as an approximation to $H(x)$. As shown in Fig. 1(a), $g(x)$ can take either a narrow and sharp or wide and flat shape; a sharper function makes its integral closer to $H(x)$, implying a more accurate gradient approximation.
@@
We aim to control the gradient-vanishing degree by adjusting the sharpness of the surrogate function for a given surrogate form. We adopt the Atan surrogate gradient (see Eq. 7) because it is smooth and has a long tail, which is favorable for stabilizing backpropagation in SNNs. From Theorem 1, we can select an appropriate sharpness parameter $\alpha$ such that the expected gradient-vanishing degree is upper-bounded by a prescribed constant:
@@
Let $g(x)=\frac{\alpha}{2\left[1+(\frac{\pi}{2}\alpha x)^{2}\right]}$ and the corresponding gradient-vanishing degree function $G(x)=\int_{-x}^{x}g(t)\,dt,x\geq 0$. Assume that $x$ follows a probability distribution $p(x)$ with finite expectation $\mathbb{E}[x]<+\infty$. Then, for any constant $A\in[0,1]$, the expected gradient-vanishing degree satisfies
$\mathbb{E}_{x\!\sim\!p(x)}[G(x)]\leq A\;\text{if}\;\alpha\leq\frac{2}{\pi\mathbb{E}[x]}\tan\!\!\left(\frac{\pi A}{2}\right)$ (9)
The detailed proof required for this section is provided in the Supplementary Materials.
@@ 4.2 Adaptive Sharpness Surrogate Gradient
$\alpha_{i,t}^{l}=\frac{2}{\pi\,\mathbb{E}[\left\lvert u_{i,t}^{l}\right\rvert]}\tan\!\left(\frac{\pi}{2}A\right).$ (10)
where $A\in[0,1]$ is a tunable constant. We replace the inequality in Eq. 9 with an equality, as we prefer $\alpha_{i,t}^{l}$ to take the maximum admissible value under the constraint $\mathbb{E}_{\left\lvert u_{i,t}^{l}\right\rvert\sim p(\left\lvert u_{i,t}^{l}\right\rvert)}\!\left[G(\left\lvert u_{i,t}^{l}\right\rvert)\right]\leq A$, ensuring that the surrogate gradient approximates $\frac{dH}{dx}$ as closely as possible.
@@
$\alpha_{i,t}^{l}(\boldsymbol{x_{1:k}})=\frac{2}{\pi\,M_{i,t}^{l}(\boldsymbol{x_{1:k}})}\tan\!\left(\frac{\pi}{2}A\right).$ (12)
@@
$\alpha_{i,t}^{l}(\boldsymbol{x_{1:k}})=\frac{2}{\pi\,\left[M_{i,t}^{l}(\boldsymbol{x_{1:k}})+\gamma D_{i,t}^{l}(\boldsymbol{x_{1:k}})\right]}\tan\!\left(\frac{\pi}{2}A\right).$ (15)
where $\beta_{1},\beta_{2}\in(0,1)$ are the momentum coefficients, $\gamma$ is the relaxation coefficient. The relaxation term is introduced to slightly reduce $\alpha_{i,t}^{l}$ by incorporating the second-order moment information of the $\left\lvert u_{i,t}^{l}\right\rvert$ distribution, thereby ensuring that the degree of gradient vanishing remains within the theoretical upper bound.
As shown in Fig. 1(b), $M_{i,t}^{l}$ and $D_{i,t}^{l}$ are initialized to $M_{0}$ and $D_{0}$. Eq. 13-Eq. 15 together constitute the adaptive update rule of ASSG, which allows each neuron to self-adjust its surrogate sharpness.
Figure 2: The distribution of the sharpness parameter $\alpha_{i,t}^{l}$. The vertical lines denote the mean and the 90th percentile (P90), respectively.
In Fig. 2, we visualize the distribution of the sharpness parameter $\alpha_{i,t}^{l}$ across different spatial-temporal dimensions in ASSG and its evolution along the optimization trajectory. The figure also marks the mean and the 90th percentile (P90) lines. As shown, $\alpha_{i,t}^{l}$ gradually develops into a unimodal, long-tailed distribution as the iterations proceed, indicating that both sharp and smooth surrogate gradient shapes coexist within the SNN. This highly spatial-temporally heterogeneous and adaptively adjusted backpropagation mechanism ensures stable gradient flow throughout the adversarial optimization trajectory.
@@ 5 Experiments
Datasets Methods Clean Attack BPTR [ding2022snn] RGA [bu2023rate] PDSG [lun2025towards] HART [hao2023threaten] STBP [wu2018spatio] ASSG

### v2

@@ 1 Introduction
Spiking Neural Networks (SNNs) have emerged as brain-inspired models that emulate the sparse, event-driven processing of biological neural systems [36, 32, 4, 35, 14]. By transmitting information through discrete spikes instead of continuous activations, SNNs achieve high energy efficiency and rich temporal dynamics, making them attractive for neuromorphic and low-power applications [30, 31, 7, 28, 23]. However, the discontinuous and binary nature of spike activations hinders the application of gradient-based adversarial attacks, making reliable robustness evaluation difficult [1].
To overcome the non-differentiability of spike activations, surrogate gradient methods approximate the spike derivative with a smooth function [32, 27]. While this enables backpropagation, it introduces approximation errors that distort gradients, especially under adversarial perturbations where accurate input gradients are essential [1, 2]. As a result, gradient-based robustness evaluation in SNNs remains unreliable.
Several studies have refined surrogate gradients [3, 15, 22, 34], yet existing methods lack input-dependent designs and theoretical analysis of gradient vanishing. Moreover, most works overlook the development of constrained optimization algorithms for adversarial attacks, which are equally crucial for reliable robustness evaluation.
@@
  • We design a new adversarial optimization algorithm with adaptive step-size, which achieves more stable and efficient convergence even when gradients are imprecise.
@@
  • Our findings reveal that the adversarial robustness of current SNNs has been significantly overestimated. Without ASSG, SNNs may even exhibit false robustness, highlighting the urgent need to rethink how to efficiently generate reliable adversarial samples for training more genuinely robust SNN models.
@@ 2.1 Gradient Approximation Methods in SNNs
Recent studies have shown that SNNs are vulnerable to adversarial perturbations [8, 20]. Surrogate gradient methods approximate the spike activation with a continuous function to enable Spatial-Temporal Backpropagation (STBP) [32]. The Potential-Dependent Surrogate Gradient (PDSG) method further incorporates the variance of membrane potential distributions into the surrogate computation [22]. However, PDSG derives the surrogate function from batch-level statistics rather than individual samples, preventing adaptive modeling of sample-specific dynamics and lacking theoretical justification for its formulation.
Another line of research estimates input gradients by differentiating the SNN loss with respect to neurons’ average firing rates. Representative methods include Backward Pass Through Rate (BPTR) [8] and Rate Gradient Approximation (RGA) [3]. The Hybrid Adversarial attack based on Rate and Temporal information (HART) [15] jointly considers rate- and time-based dependencies, but its derivation is limited to a specific neuronal model and lacks theoretical validation for generalization to broader neuron types.
@@ 2.2 Adversarial Defense Methods for SNNs
Early studies adapted adversarial training for SNNs, such as adversarial input augmentation and min–max optimization [24]. Later approaches, including Regularized Adversarial Training (RAT), constrain the Lipschitz constant of each layer via weight regularization to stabilize learning [8], while Sparsity Regularization (SR) introduces a sparse gradient penalty into the loss to enhance robustness when combined with adversarial training [21].
More recent works have strengthened SNN adversarial defense from an architectural perspective [10, 33]. For instance, synaptic gating mechanisms dynamically regulate spike transmission to suppress adversarial noise propagation [10], and frequency-domain encoding schemes improve robustness by representing spiking information in more stable spectral domains [33].
@@ 3.1 Adversarial Attack and Training
Adversarial attacks can be viewed as a constrained optimization problem in the input space of a neural network [25, 5, 24, 6, 11]. For a classification model $\boldsymbol{f}_{\boldsymbol{\theta}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$, let $\boldsymbol{x}\in\mathbb{R}^{d}$ denote a benign input sample with the ground-truth label $y$. The objective of an untargeted adversarial attack is to find a perturbed input $\boldsymbol{x^{\prime}}$ that maximizes the classification loss $\mathcal{L}$ while remaining within a small $L_{p}$-norm ball around the original input:
@@
here, $\varepsilon$ defines the perturbation bound that controls the perceptual similarity between $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$. In practice, gradient-based optimization methods such as Projected Gradient Descent (PGD) and Auto-PGD (APGD) are commonly used to approximate the solution of this problem under various norm constraints (e.g. $L_{\infty}$, $L_{2}$[24, 6].
Adversarial Training (AT) is one of the most fundamental methods for improving a network’s adversarial robustness, and it has also been widely adopted in SNN-related studies [8, 9, 10, 33, 21, 26, 18]. The objective of AT can be formulated as a min–max optimization problem:
@@
where $\mathcal{D}$ denotes the data distribution. TRadeoff-inspired Adversarial DEfense via Surrogate-loss minimization (TRADES) is one of the most influential variants of AT [37, 29], which is also employed as a training framework for developing robust SNNs in this work.
@@ 3.2 Spiking Neurons and Surrogate Gradients
To evaluate the generality of the proposed attack, we employ several neuron models in addition to the standard LIF neuron, including its variant LIF-2, the Integrate-and-Fire (IF) neuron, and the Parallel Spiking Neuron (PSN) [14]. Detailed configurations of these models are provided in the Appendix Sec. S2.
@@
where $\alpha$ controls the sharpness of the surrogate gradient. In the Appendix Sec. S1, we provide additional examples of surrogate gradient functions.
@@ 4 Methods
Figure 1: (a). Curves of the surrogate gradient function $g(x)$ and its integral under different sharpness parameters $\alpha$. (b). Heterogeneous evolution of $\left|u_{i,t}^{l}\right|$ and $\alpha_{i,t}^{l}$ across different inputs at a fixed $(i,t,l)$ during adversarial optimization. (c) spatial-temporal heterogeneity of a group of $\alpha_{i,t}^{l}$ across different inputs. The Pearson correlation coefficients of the input-dependent $\alpha_{i,t}^{l}$ are also shown in the figure.
@@ 4.1 Theoretical Analysis on Gradient Vanishing
Surrogate gradient is employed to approximate a non-zero and smooth derivative of $H(x)$ for training or adversarial attack. We define the surrogate gradient as the following general class of functions:
@@
Unlike the definition given by Mukhoty et al. [27], we impose the normalization constraint $\int_{-\infty}^{\infty}g(x)\,dx=1$. This property guarantees that the integral of $g(x)$ behaves as an approximation to $H(x)$. As shown in Fig. 1(a), $g(x)$ can take either a narrow and sharp or wide and flat shape; a sharper function makes its integral closer to $H(x)$, implying a more accurate gradient approximation.
@@
We aim to control the gradient-vanishing degree by adjusting the sharpness of the surrogate function. Without loss of generality, we express the surrogate function in the form $\alpha k(\alpha x)$, whose primitive function can be written as $\int_{-\infty}^{x}\alpha k(\alpha x)dx=K(\alpha x)$. The parameter $\alpha$ in this surrogate function can be used to control its sharpness. As shown in Fig. 1(a), a larger $\alpha$ leads to a sharper surrogate function, and its primitive function becomes closer to $H(x)$, indicating more accurate gradient estimation.
From Theorem 1, we can select an appropriate sharpness parameter $\alpha$ such that the expected gradient-vanishing degree is upper-bounded by a prescribed constant:
@@
Let $g(x)=\alpha k(\alpha x)$ a surrogate function, and the corresponding gradient-vanishing degree function $\int_{-x}^{x}\alpha k(\alpha t)\,dt=\int_{-\alpha x}^{\alpha x}k(\alpha t)\,d\alpha t=G(\alpha x),x\geq 0$. Assume that $x$ follows a probability distribution $p(x)$ with finite expectation $\mathbb{E}[x]<+\infty$. Then, for any constant $A\in[0,1]$, the expected gradient-vanishing degree satisfies
$\mathbb{E}_{x\!\sim\!p(x)}[G(\alpha x)]\leq A\;\text{if}\;\alpha\leq\frac{G^{-1}(A)}{\mathbb{E}[x]}$ (9)
The detailed proof required for this section is provided in the Appendix Sec. S1.
@@ 4.2 Adaptive Sharpness Surrogate Gradient
$\alpha_{i,t}^{l}=\frac{\omega}{\mathbb{E}[\left\lvert u_{i,t}^{l}\right\rvert]}$ (10)
where $\omega=G^{-1}(A)$. We replace the inequality in Eq. 9 with an equality, as we prefer $\alpha_{i,t}^{l}$ to take the maximum admissible value under the constraint $\mathbb{E}_{\left\lvert u_{i,t}^{l}\right\rvert\sim p(\left\lvert u_{i,t}^{l}\right\rvert)}\!\left[G(\left\lvert u_{i,t}^{l}\right\rvert)\right]\leq A$, ensuring that the surrogate gradient approximates $\frac{dH}{dx}$ as closely as possible. $\omega$ is tunable to adjust the upper bound of gradient-vanishing degree. To obtain the optimal gradient accuracy-vanishing trade-off for attack, we select the best $\omega$ via ternary search.
For different surrogate functions, the valid range of $\omega$ may vary. However, we uniformly set its range to $0,+\infty)$. In the Appendix Sec. S1, we provide a detailed explanation of this range and give examples of the specific computation of $\omega$ for various surrogate functions.
@@
$\alpha_{i,t}^{l}(\boldsymbol{x_{1:k}})=\frac{\omega}{M_{i,t}^{l}(\boldsymbol{x_{1:k}})}.$ (12)
@@
$\alpha_{i,t}^{l}(\boldsymbol{x_{1:k}})=\frac{\omega}{M_{i,t}^{l}(\boldsymbol{x_{1:k}})+\gamma D_{i,t}^{l}(\boldsymbol{x_{1:k}})}.$ (15)
where $\beta_{1},\beta_{2}\in(0,1)$ are the momentum coefficients, $\gamma$ is the relaxation coefficient. The relaxation term is introduced to slightly reduce $\alpha_{i,t}^{l}$ by incorporating the second-order moment information of the $\left\lvert u_{i,t}^{l}\right\rvert$ distribution, thereby ensuring that the gradient-vanishing degree remains within the theoretical upper bound.
$M_{i,t}^{l}$ and $D_{i,t}^{l}$ are initialized to $M_{0}$ and $D_{0}$. Eq. 13-Eq. 15 together constitute the adaptive update rule of ASSG, which allows each neuron to self-adjust its surrogate sharpness.
Figure 2: The distribution of the sharpness parameter $\alpha_{i,t}^{l}$. The vertical lines denote the mean and the 90th percentile (P90).
In Fig. 1(b), We visualize, at a fixed spatial-temporal location $(i,t,l)$, the dynamics of $\left\lvert u_{i,t}^{l}\right\rvert$ for different inputs during the adversarial attack process, together with the input-adaptive $\alpha_{i,t}^{l}$ computed by ASSG. It can be observed that the closer the distribution of $\left|u_{i,t}^{l}\right|$ is to 0, the larger the $\alpha_{i,t}^{l}$ evolved by ASSG.
In Fig. 1(c), we visualize the distribution differences of $\alpha_{i,t}^{l}$ at a group of spatial-temporal locations $(i,t,l)$ for different inputs at the 75th iteration of the adversarial attack. The figure shows that the pairwise Pearson correlation coefficients among the three groups of $\alpha_{i,t}^{l}$ are all very low, further demonstrating that different inputs lead to highly heterogeneous spatial-temporal distributions of $\alpha_{i,t}^{l}$. This indicates that ASSG possesses strong input-adaptive capability across both spatial and temporal dimensions.
In Fig. 2, we visualize the distribution of the sharpness parameter $\alpha_{i,t}^{l}$ across different spatial-temporal dimensions in ASSG and its evolution along the optimization trajectory. As shown, $\alpha_{i,t}^{l}$ gradually develops into a more dispersed unimodal, long-tailed distribution as the iterations proceed, indicating that both sharp and smooth surrogate gradient shapes coexist within the SNN.
In Fig. 3, we perform backpropagation in the SNN using the Atan surrogate gradient with different fixed $\alpha$ values, as well as with ASSG that adaptively computes $\alpha_{i,t}^{l}$ using $A=0.87$, and visualize the distribution of gradient-vanishing degree $G(x)$ for neurons across all spatial-temporal dimensions. When fixed $\alpha$ is used, the distribution of $G(x)$ in the SNN is more dispersed (with higher entropy). As $\alpha$ increases, the distribution shifts toward 1. In contrast, ASSG effectively concentrates the gradient vanishing degrees in the SNN around the theoretically upper bound $A$ (with the distribution peak at 0.8669).
Figure 3: The distribution of the degree of gradient vanishing $G(x)$ in the SNN using the Atan surrogate gradient with different fixed $\alpha$ and with ASSG ($A=0.87$)
In summary, Fig. 1(b)(c) demonstrates the input-dependent property of ASSG, while Fig. 2 shows that, for each input, ASSG evolves a highly spatial-temporally heterogeneous distribution of $\alpha_{i,t}^{l}$. Furthermore, ASSG enables the surrogate gradients of every neuron to be as accurate as possible (as reflected by the long-tailed distribution in Fig. 2), while constraining the gradient-vanishing degree in the SNN to be tightly concentrated within a specified upper bound (Fig. 3). Consequently, by adjusting this upper bound, we can allow ASSG to perform backpropagation under different gradient accuracy-vanishing trade-offs, thereby achieving stronger attack performance.
@@ 5 Experiments
Datasets Methods Clean Attack BPTR [8] RGA [3] PDSG [22] HART [15] STBP [32] ASSG
@@
Table 2: ASR on different surrogate gradient functions.
Method Atan Rect Tri Sig Gauss
CIFAR-10
fixed $\alpha$ 84.04 77.90 81.07 81.89 81.29
ASSG 88.44 86.82 89.53 89.30 89.46
CIFAR-100
fixed $\alpha$ 91.40 90.02 91.08 91.22 91.16
ASSG 93.19 92.31 93.41 93.35 93.42